# Prompt Guiding Multi-Scale Adaptive Sparse Representation-driven Network for Low-Dose CT MAR

### v1

@@ Abstract
Low-dose CT (LDCT) is capable of reducing X-ray radiation exposure, but it will potentially degrade image quality, even yields metal artifacts at the case of metallic implants. For simultaneous LDCT reconstruction and metal artifact reduction (LDMAR), existing deep learning-based efforts face two main limitations: i) the network design neglects multi-scale and within-scale information; ii) training a distinct model for each dose necessitates significant storage space for multiple doses. To fill these gaps, we propose a prompt guiding multi-scale adaptive sparse representation-driven network, abbreviated as PMSRNet, for LDMAR task. Specifically, we construct PMSRNet inspired from multi-scale sparsifying frames, and it can simultaneously employ within-scale characteristics and cross-scale complementarity owing to an elaborated prompt guiding scale-adaptive threshold generator (PSATG) and a built multi-scale coefficient fusion module (MSFuM). The PSATG can adaptively capture multiple contextual information to generate more faithful thresholds, achieved by fusing features from local, regional, and global levels. Furthermore, we elaborate a model interpretable dual domain LDMAR framework called PDuMSRNet, and train single model with a prompt guiding strategy for multiple dose levels. We build a prompt guiding module, whose input contains dose level, metal mask and input instance, to provide various guiding information, allowing a single model to accommodate various CT dose settings. Extensive experiments at various dose levels demonstrate that the proposed methods outperform the state-of-the-art LDMAR methods.
Index Terms:
Low-dose computed tomography, Metal artifact reduction, Multi-scale sparse representation, Prompt guiding, Interpretability.
I Introduction
X-ray computed tomography (CT) has been extensively utilized for medical diagnosis. However, the X-ray radiation increases the risk of cancer to patients when aggregated exposure surpasses the specific threshold (Bosch de Basea Gomez et al., 2023). Typically, a reduction in radiation dose is accomplished by managing the electrical current or voltage of the X-ray tube. Consequently, low-dose CT (LDCT) is extensively utilized to mitigate X-ray radiation risks. However, the decrease of the X-ray flux level leads to a sinogram (X-ray projection) corrupted by Poisson noise, subsequently yielding the reconstruction corrupted by noise. Additionally, when metallic implants are present within patients, the quality of LDCT images will be further reduced by metal artifacts (Zhou et al., 2022). Therefore, how to simultaneously perform low-dose CT denoising and metal artifact reduction (LDMAR) is a challenge.
With the continuous advancement of deep learning (DL) in medical image processing, various DL-based approaches using deep neural networks (DNNs) have been introduced into the LDCT denoising and full-dose CT (FDCT) MAR. Existing DL-based LDCT denoising or FDCT MAR methods can be primarily categorized into three classes: sinogram domain (Ma et al., 2021; Ghani and Karl, 2020), image domain (Wang et al., 2023a; Liang et al., 2020; Yang et al., 2018; Liao et al., 2020; Li et al., 2023a), and dual domain (Ge et al., 2022; Wu et al., 2023; Wang et al., 2021) methods. Specifically, sinogram domain methods are designed to correct the sinogram and subsequently apply the filtered back-projection (FBP) operation for CT image reconstruction. Image domain methods build DNNs to learn mapping between clean CT images and degraded CT images. In contrast, dual domain methods integrate knowledge from both the sinogram domain and the image domain to construct dual domain networks, achieving significant success. Although aforementioned methods have been developed for LDCT denoising and FDCT MAR tasks, applying previous techniques directly to the LDMAR task produces unsatisfactory reconstructions. To address this issue, [42] first proposed DuDoUFNet, a dual domain under-to-fully-complete progressive restoration network for LDMAR. Nevertheless, DuDoUFNet neglects both multi-scale and within-scale information, limiting its performance, and it also suffers from inadequate model interpretability. In addition, the aforementioned methods suffer from large storage costs at the case of multiple dose levels since they often train each model for the specific setting. This “one-model-for-one setting” approach not only increases storage overhead but also produces inflexible models tailored to single settings, limiting their adaptability in clinical applications. To overcome these bottlenecks, we propose a prompt guiding multi-scale adaptive sparse representation-driven network called PMSRNet which can train a single model for multiple doses, and incorporate it into a dual domain framework to build a model interpretable network called PDuMSRNet. The main contributions can be summarized as follows:
(1) We come up with a prompt guiding multi-scale adaptive sparse representation-driven network dubbed as PMSRNet. It employs multi-scale sparsifying frames to establish a model-driven multi-scale adaptive network that incorporates both within-scale characteristics and cross-scale complementarity into the overall network. Serving as a plug-and-play module, PMSRNet can be easily generalized to address various image restoration tasks. Specially, we leverage the dose map, metal mask and input instance as prompts, and embedding them into a deep unfolded framework for solving LDMAR task.
(2) We build a prompt guiding scale-adaptive threshold generator (PSATG) capable of adaptively extracting within-scale features for different scales to generate faithful thresholds during the shrinking procedure. The PSATG is composed of a shallow feature extraction module, a deep feature extraction module, a feature selection module, a threshold generating module, and a prompt guiding module. By using a prompt guiding module, the guiding information can be injected into the procedure of generating thresholds under the case of multiple dose levels.
(3) We propose a model interpretable dual domain LDMAR framework. Specifically, we formulate a sparsifying transform-based dual domain optimization model, and solve this optimization problem by using proximal gradient methods. Subsequently, we unfold this iteration algorithm into a deep neural network whose image domain subnetwork is the same architecture as that of PMSRNet, and the sinogram domain subnetwork consists of convolutional residual networks. The comprehensive experiments on the LDMAR task show that the proposed method can produce better results compared to SOTA methods.
II Related Work
II-A LDCT Denoising
For LDCT denoising, numerous algorithms have been developed to improve the quality of LDCT images. Traditional methods can be summarized into three main groups, namely sinogram domain filtering (Balda et al., 2012), iterative reconstruction (Lee et al., 2017; Liu et al., 2017), and image post-processing (Chen et al., 2013). The sinogram domain filtering methods employ specific filters to directly suppress noise on raw projection data, potentially causing edge blurring and loss of structural information. The iterative reconstruction methods aim to minimize a unified objective function. However, these methods require raw projection data. In contrast, the image post-processing methods can directly process LDCT images, making them easily integrated into the CT machine system.
Nonetheless, it is challenging for traditional hand-crafted priors to characterize all LDCT images. Recently, DL-based methods show promising reconstruction quality for LDCT denoising due to their powerful image distribution modeling capabilities. Some studies initially developed sinogram domain networks to correct sinograms, followed by using FBP to reconstruct CT images. [8] adopted convolution residual networks to learn mapping between low-dose and full-dose sinograms. [20] proposed an attention residual dense network address the low-dose sinogram denoising problem. Minor errors in sinograms may introduce new noise and artifacts into the reconstructed images. In contrast, image domain methods directly perform on LDCT images. [4] designed a residual encoder-decoder network to address the LDCT denoising issue. [30] proposed a transformer-based network called CTformer, which incorporates dilated and cyclic shifts to comprehensively capture interactions. [15] proposed an densely connected convolutional neural network with a trainable Sobel convolution design. Since noise stems from the back-projection of corrupted projection data, modeling the noise distribution is challenging. To address the limitation of single domain networks, dual domain methods leverage information from sinogram and image domains for LDCT denoising. [38] proposed a dual domain network based on a 3D residual convolution network. [7] proposed a parallel network to fully and collaboratively utilized dual domain information. [6] proposed an attention–encoding integrated generative adversarial network to achieve efficient full-dose reconstruction. However, these LDCT algorithms do not account for the existence of metallic implants within patients. Therefore, directly applying them for LDMAR may lead to suboptimal reconstruction performance.
II-B FDCT MAR
For FDCT MAR, numerous methods have been proposed and have achieved outstanding performance. Physical effects correction methods model the physical effects of imaging and directly corrects metal-corrupted sinogram regions (Park et al., 2016). However, when high-atomic number metals are present, the metal trace region can be so severely damaged that the above correction methods fail to produce satisfactory results (Zhang and Yu, 2018). Sinogram completion methods utilize various estimation techniques to fill the metal trace region within sinogram (Meyer et al., 2009). As it is difficult to generate perfect estimation, these estimation methods often introduce secondary artifacts. To cope with this problem, iterative reconstruction methods (Chang et al., 2019; Mehranian et al., 2013) are used to reconstruct high-quality CT images from the observed data, but these methods require suitable hand-crafted regularization terms.
Figure 1: Overview of our prompt guiding multi-scale adaptive sparse representation-driven network named PMSRNet (top), our prompt guiding scale-adaptive threshold generator termed as PSATG (middle) and multi-scale coefficient fusion module called MSFuM (bottom). The PMSRNet contains three components: multi-scale sparsifying frames, PSATG and MSFuM. The PSATG contains five parts: 1) The shallow feature extraction module relies on a convolution layer; 2) The deep feature extraction module models images at local, regional, and global levels by parallelizing convolutional neural networks, window attention, and anchored stripe attention; 3) The feature selection module utilizes channel attention mechanism to adaptively reweight information from different ranges; 4) The threshold generating module enhances the performance of PMSRNet by generating faithful thresholds using information from the feature selection module; 5) The prompt guiding module utilizes prompt information, including LDCT images with metallic implants, metal masks, and dose maps, to guide the threshold generating process. The MSFuM which is used to fuse low-resolution and high-resolution images mainly contains a cross attention and a feed-forward network.
With the widespread application of DNNs in medical image processing, numerous DL-based methods have been developed for MAR. [9] employed generative adversarial network to correct metal-corrupted sinogram. Since sinogram domain methods are difficult to produce perfect estimation of the metal trace region, secondary artifacts will inevitably be introduced to the reconstructed image. Image domain methods regard the MAR task as image restoration. [41] utilized DNNs to generate the prior image that assists in correcting the metal-corrupted regions of the sinogram. [16] introduced an unsupervised artifact disentanglement network that produces artifact-reduced images using unpaired training data. The primary limitation of image domain methods is the neglect of the consistency constraint. To address the issue of single domain networks, dual domain methods that exploit dual domain knowledge have been developed these years. The seminal work termed as DuDoNet was proposed by [17], and a Radon inversion layer was designed to bridge between the sinogram domain and the image domain network. [33] proposed a parallel dual domain network by modifying the fixed priority mode of the previous dual domain network. [31] proposed a model-driven dual domain network, integrating imaging geometrical model constraints into the mutual learning process of sinogram and image domain information. [32] proposed a deep unfolding dual domain network by unfolding the iterative algorithm for MAR task. Although the dual domain network mentioned above have achieved high-quality CT images, the lack of domain knowledge during training remains a significant limitation in improving the MAR performance. [25] proposed a method that leverages sinogram domain, image domain, artifact domain, and coding domain knowledge to construct a model-driven network for MAR. [29] proposed a novel dual domain network that couples model- and data-driven networks for MAR. Although these networks have achieved excellent performance in MAR tasks, they are designed for full-dose CT and do not address the LDMAR issue. Furthermore, the design of these networks has not considered the incorporation of multi-scale and within-scale information. This leaves room for further improvement.
III Prompt Guiding Multi-Scale Adaptive Sparse Representation-Driven Network
III-A Architecture Principles
Multi-scale features and cross-scale complementarity are important for solving image inverse problems. [10] proposed MSANet that effectively integrates the within-scale characteristics and cross-scale complementarity for image denoising. Inspired by this, we propose a prompt guiding multi-scale sparse representation-driven network for LDMAR. In PMSRNet, we employ a learnable sparsifying frame to filter the input data for noise removal. More precisely, this process can be expressed as
$\textbf{\emph{x}}={\textbf{\emph{W}}}^{\mathrm{T}}S_{\varepsilon}({\textbf{\emph{W}}}\tilde{\textbf{\emph{x}}})$ (1)
where $\tilde{\textbf{\emph{x}}}$ represents the input LDCT image with metallic implants, x represents the filtered CT image, W represents the sparsifying frame, and $S_{\varepsilon}(\cdot)$ is the soft threshold function defined as $soft_{\varepsilon}(\mu)=sign(\mu)max(|\mu|-\varepsilon,0)$. To enhance the representation ability of the single-scale sparsifying frame, we extend it to multi-scale sparsifying frames. Specifically, the multi-scale sparsifying frames are sequential structure where information flows between different scales while remaining unchanged within each scale. To obtain multi-scale features, we set the filter sizes in the sparsifying frame network to 5$\times$5, 7$\times$7, 9$\times$9, and 11$\times$11, with corresponding output channel numbers of 25, 49, 81, and 121, respectively. The threshold in the soft threshold function plays a crucial role in the sparsifying frame network. Therefore, we construct a prompt guiding scale-adaptive threshold generator to generate faithful thresholds.
III-B Prompt Guiding Scale-Adaptive Threshold Generator
Both high-resolution and low-resolution layers require contextual information. When it comes to high-resolution features, reducing their resolution can result in the loss of critical image details. Conversely, if the resolution of low-resolution features is insufficient, it may negatively affect the structural consistency of the reconstructed image. To achieve complementarity of within-scale information, we propose a prompt guiding scale-adaptive threshold generator (PSATG) that can fuse the prompting features and various features from local, regional, and global levels. Specifically, it consists of a shallow feature extraction module, a deep feature extraction module, a feature selection module, a threshold generating module, and a prompt guiding module. The shallow feature extraction module is based on a convolution layer. The deep feature extraction module contains a local information extraction branch (LIEB), a regional information extraction branch (RIEB), and a global information extraction branch (GIEB). LIEB adopts a convolutional residual structure to extract local information. The self-attention branches RIEB and GIEB are built on window attention (Liu et al., 2022) and anchored stripe attention (Li et al., 2023b) to extract regional and global information, respectively. The feature maps are independently processed within three branches and later combined by concatenating them along the channel dimension. The prompt guiding module utilizes prompt information to guide the threshold generating process. More precisely, the initial input feature map undergoes processing through the shallow feature extraction module to generate ${\hat{\bm{x}}}$. Next, we perform the three branches on ${\hat{\bm{x}}}$ to extract local, regional, and global information, respectively. The feature selection module reweights channels corresponding to the three branches, selecting information more conducive to image reconstruction. Finally, the reweighted information is input into the prompt guiding threshold generation module for fusion.
Recently, prompt learning is popular for addressing the all-in-one image restoration (IR) tasks (Potlapalli et al., 2023). This approach not only facilitates generalization across various degradation types and levels, but also consistently achieves SOTA results in IR tasks. However, it often requires various datasets to effectively capture the discriminant features of several IR tasks. For LDMAR task, it is difficult to build various CT datasets. To address this issue, and inspired from prompt learning, we utilize the explicit prompts to guide the threshold generating process. By doing so, the discriminant information for different dose levels can be used to guide the reconstruction network. In this paper, we do not focus on the learning of discriminative information, but rather on improving the overall LDMAR performance. As illustrated in Fig. 1, we take the LDCT images with metallic implants, the metal masks and the dose maps as prompts, and perform a channel-wise concatenation operation on them to obtain the three-channel image P. A prompt guiding sub-module represented by ${F_{ext}({\bullet})}$ is employed to extract features from the prompt P. Subsequently, these features are processed through fully connected layers ${\rm FC({\bullet})}$ to achieve the appropriate shape (1$\times$×channel or dimension) required for the corresponding module’s outputs. Finally, the output features are multiplied by weight k and added by bias b. Mathematically, the process of PSATG can be formulated as
$\textbf{\emph{k}}_{1}={\rm FC}_{1}(F_{\rm ext}(\textbf{P})),\textbf{\emph{k}}_{2}={\rm FC}_{2}(F_{\rm ext}(\textbf{P})),$ (2)
$\textbf{\emph{b}}_{1}={\rm FC}_{3}(F_{\rm ext}(\textbf{P})),\textbf{\emph{b}}_{2}={\rm FC}_{4}(F_{\rm ext}(\textbf{P})),$ (3)
$\textbf{\emph{x}}_{\rm in}={\rm FSM}(\textbf{\emph{k}}_{1}[{\rm LIEB}(\textbf{\emph{x}}_{\rm in}),{\rm RIEB}(\textbf{\emph{x}}_{\rm in}),{\rm GIEB}(\textbf{\emph{x}}_{\rm in})]+\textbf{\emph{b}}_{1}),$ (4)
and
$\textbf{\emph{x}}_{\rm out}={\rm Conv}({\rm ReLU}({\rm Conv}({\textbf{\emph{k}}_{2}}\textbf{\emph{x}}_{\rm in}+{\textbf{\emph{b}}_{2}})))$ (5)
where FSM($\cdot$) denotes the feature selection module, and [$\textbf{\emph{x}}_{1}$, $\textbf{\emph{x}}_{2}$, $\textbf{\emph{x}}_{3}$] denotes the channel-wise concatenation of $\textbf{\emph{x}}_{1}$, $\textbf{\emph{x}}_{2}$, and $\textbf{\emph{x}}_{3}$.
@@ III-B Prompt Guiding Scale-Adaptive Threshold Generator
III-C Multi-Scale Coefficient Fusion Module
$\tilde{\textbf{\emph{Q}}}&=\rm R(Dconv(Conv(\textbf{\emph{Q}}))\\ \tilde{\textbf{\emph{K}}}&=\rm R(Dconv(Conv(\textbf{\emph{K}}))\\ \tilde{\textbf{\emph{V}}}&=\rm R(Dconv(Conv(\textbf{\emph{V}}))\end{aligned}\right.$ (6)
$\textbf{\emph{C}}=\rm SoftMax(\tilde{\textbf{\emph{K}}}^{T}\tilde{\textbf{\emph{Q}}}).$ (7)
where $\rm{SoftMax(}\bullet{)}$ is the softmax function applied to re-weight the matrix multiplication. Next, we employ the reshape and convolution operations to further extract features. In summary, cross attention of MSFuM can be expressed as
$Attention(\textbf{\emph{Q}},\textbf{\emph{K}},\textbf{\emph{V}})=\rm Conv(R(\tilde{\textbf{\emph{V}}}\textbf{\emph{C}})).$ C ) ) . (8)
IV The proposed PMSRNet for Dual Domain LDMAR Network
IV-A Model Design and Optimization

### v3

@@ Abstract
Low-dose CT (LDCT) is capable of reducing X-ray radiation exposure, but it will potentially degrade image quality, even yields metal artifacts at the case of metallic implants. For simultaneous LDCT reconstruction and metal artifact reduction (LDMAR), existing deep learning-based efforts face two main limitations: i) the network design neglects multi-scale and within-scale information; ii) training a distinct model for each dose necessitates significant storage space for multiple doses. To fill these gaps, we propose a prompt guiding multi-scale adaptive sparse representation-driven network, abbreviated as PMSRNet, for LDMAR task. Specifically, we construct PMSRNet inspired from multi-scale sparsifying frames, and it can simultaneously employ within-scale characteristics and cross-scale complementarity owing to an elaborated prompt guiding scale-adaptive threshold generator (PSATG) and a built multi-scale coefficient fusion module (MSFuM). The PSATG can adaptively capture multiple contextual information to generate more faithful thresholds, achieved by fusing features from local, regional, and global levels. Furthermore, we elaborate a model interpretable dual domain LDMAR framework called PDuMSRNet, and train single model with a prompt guiding strategy for multiple dose levels. We build a prompt guiding module, whose input contains dose level, metal mask and input instance, to provide various guiding information, allowing a single model to accommodate various CT dose settings. Extensive experiments at various dose levels demonstrate that the proposed methods outperform the state-of-the-art LDMAR methods. Code is available in https://github.com/shibaoshun/PDuMSRNet.git
keywords:
2024 MSC:
41A05
41A10
65D05
@@
Low-dose computed tomography
Metal artifact reduction
Multi-scale sparse representation
Prompt guiding
Interpretability
journal: Medical Image Analysis
1 Introduction
X-ray computed tomography (CT) has been extensively utilized for medical diagnosis. However, the X-ray radiation increases the risk of cancer to patients when aggregated exposure surpasses the specific threshold (9762023, 9762023). Typically, a reduction in radiation dose is accomplished by managing the electrical current or voltage of the X-ray tube. Consequently, low-dose CT (LDCT) is extensively utilized to mitigate X-ray radiation risks. However, the decrease of the X-ray flux level leads to a sinogram (X-ray projection) corrupted by Poisson noise, subsequently yielding the reconstruction corrupted by noise. Additionally, when metallic implants are present within patients, the quality of LDCT images will be further reduced by metal artifacts (9825711, 9825711). Therefore, how to simultaneously perform low-dose CT denoising and metal artifact reduction (LDMAR) is a challenge.
With the continuous advancement of deep learning (DL) in medical image processing, various DL-based approaches using deep neural networks (DNNs) have been introduced into the LDCT denoising and full-dose CT (FDCT) MAR. Existing DL-based LDCT denoising or FDCT MAR methods can be primarily categorized into three classes: sinogram domain ( 8815915, 8815915; 9439999, 9439999), image domain (8340157, 8340157; 9320928, 9320928; 8788607, 8788607; 9362125, 9362125; 9852485, 9852485; Wang2022DomainAD, Wang2022DomainAD; Wang_2023, Wang_2023; 9928347, 9928347; mia2024ldct, mia2024ldct), and dual domain ( 10.1007/978-3-031-16446-0_71, 10.1007/978-3-031-16446-0_71; 10.1007/978-3-030-87231-1_11, 10.1007/978-3-030-87231-1_11; 10309982, 10309982; mia2025ldct1, mia2025ldct1; mia2025ldct3, mia2025ldct3) methods. Specifically, sinogram domain methods are designed to correct the sinogram and subsequently apply the filtered back-projection (FBP) operation for CT image reconstruction. Image domain methods build DNNs to learn mapping between clean CT images and degraded CT images. In contrast, dual domain methods integrate knowledge from both the sinogram domain and the image domain to construct dual domain networks, achieving significant success. Although aforementioned methods have been developed for LDCT denoising and FDCT MAR tasks, applying previous techniques directly to the LDMAR task produces unsatisfactory reconstructions. To address this issue, [9825711] first proposed DuDoUFNet, a dual domain under-to-fully-complete progressive restoration network for LDMAR. Nevertheless, DuDoUFNet neglects both multi-scale and within-scale information, limiting its performance, and it also suffers from inadequate model interpretability. In addition, the aforementioned methods suffer from large storage costs at the case of multiple dose levels since they often train each model for the specific setting. This “one-model-for-one setting" approach not only increases storage overhead but also produces inflexible models tailored to single settings, limiting their adaptability in clinical applications. To overcome these bottlenecks, we propose a prompt guiding multi-scale adaptive sparse representation-driven network called PMSRNet which can train a single model for multiple doses, and incorporate it into a dual domain framework to build a model interpretable network called PDuMSRNet. The main contributions can be summarized as follows:
  • 1.
    We come up with a prompt guiding multi-scale adaptive sparse representation-driven network dubbed as PMSRNet. It employs multi-scale sparsifying frames to establish a model-driven multi-scale adaptive network that incorporates both within-scale characteristics and cross-scale complementarity into the overall network. Specially, we leverage the dose map, metal mask and input instance as prompts, embedding them into a deep unfolded framework for solving the LDMAR task. PMSRNet can serve as a prior network embedded into deep unfolding networks [10.1007/978-3-030-87231-1_11] for other image reconstruction tasks.
  • 2.
    We build a prompt guiding scale-adaptive threshold generator (PSATG) capable of adaptively extracting within-scale features for different scales to generate faithful thresholds during the shrinking procedure. The PSATG is composed of a shallow feature extraction module, a deep feature extraction module, a feature selection module, a threshold generating module, and a prompt guiding module. By using a prompt guiding module, the guiding information can be injected into the procedure of generating thresholds under the case of multiple dose levels.
  • 3.
    We propose a model interpretable dual domain LDMAR framework. Specifically, we formulate a sparsifying transform-based dual domain optimization model, and solve this optimization problem by using proximal gradient methods. Subsequently, we unfold this iteration algorithm into a deep neural network whose image domain subnetwork is the same architecture as that of PMSRNet, and the sinogram domain subnetwork consists of convolutional residual networks. The comprehensive experiments on the LDMAR task show that the proposed method can produce better results compared to SOTA methods.
2 Related Work
2.1 LDCT Denoising
For LDCT denoising, numerous algorithms have been developed to improve the quality of LDCT images. Traditional methods can be summarized into three main groups, namely sinogram domain filtering (6151161, 6151161), iterative reconstruction (8080249, 8080249; 8081818, 8081818), and image post-processing (6610425, 6610425). The sinogram domain filtering methods employ specific filters to directly suppress noise on raw projection data, potentially causing edge blurring and loss of structural information. The iterative reconstruction methods aim to minimize a unified objective function. However, these methods require raw projection data. In contrast, the image post-processing methods can directly process LDCT images, making them easily integrated into the CT machine system.
Nonetheless, it is challenging for traditional hand-crafted priors to characterize all LDCT images. Recently, DL-based methods show promising reconstruction quality for LDCT denoising due to their powerful image distribution modeling capabilities. Some studies initially developed sinogram domain networks to correct sinograms, followed by using FBP to reconstruct CT images. [5401721] adopted convolution residual networks to learn mapping between low-dose and full-dose sinograms. [9439999] proposed an attention residual dense network address the low-dose sinogram denoising problem. Minor errors in sinograms may introduce new noise and artifacts into the reconstructed images. In contrast, image domain methods directly perform on LDCT images. [7947200] proposed a residual encoder-decoder network to address the LDCT denoising issue. [9320928] designed an densely connected convolutional neural network with a trainable Sobel convolution design. [9362125] proposed a deep CNN denoising approach leveraging anatomical prior information and anatomical attention mechanisms for LDCT. [Wang2022DomainAD] introduced a noise domain adaptive denoising network, which leveraged noise estimation and transfer learning. [9852485] designed an unsupervised method with noise-characteristics modeling framework. [Wang_2023] introduced a transformer-based network called CTformer, which incorporates dilated and cyclic shifts to comprehensively capture interactions. [mia2024ldct] designed an unsupervised framework for LDCT task, demonstrating considerable potential for real-world CT imaging tasks lacking paired training data. Since noise stems from the back-projection of corrupted projection data, modeling the noise distribution is challenging. To address the limitation of single domain networks, dual domain methods leverage information from sinogram and image domains for LDCT denoising. [8718010] introduced a dual domain network based on a 3D residual convolution network. [10.1007/978-3-031-16446-0_71] designed a parallel network to fully and collaboratively utilized dual domain information. [FU2023102787] proposed a generative adversarial network with attention encoding to achieve efficient full-dose reconstruction. [mia2025ldct3] introduced a dual-domain joint optimization LDCT imaging framework for noisy sparse-view LDCT image reconstruction. [mia2025ldct1] proposed a noise-inspired diffusion model for generalizable LDCT reconstruction, which requires only normal-dose data for training and can be effectively extended to various unseen dose levels. However, these LDCT algorithms do not account for the existence of metallic implants within patients. Therefore, directly applying them for LDMAR may lead to suboptimal reconstruction performance.
2.2 FDCT MAR
For FDCT MAR, numerous methods have been proposed and have achieved outstanding performance. Physical effects correction methods model the physical effects of imaging and directly corrects metal-corrupted sinogram regions (7268881, 7268881). However, when high-atomic number metals are present, the metal trace region can be so severely damaged that the above correction methods fail to produce satisfactory results (8331163, 8331163). Sinogram completion methods utilize various estimation techniques to fill the metal trace region within sinogram (5401722, 5401722). As it is difficult to generate perfect estimation, these estimation methods often introduce secondary artifacts. To cope with this problem, iterative reconstruction methods (8576532, 8576532; 6522894, 6522894) are used to reconstruct high-quality CT images from the observed data, but these methods require suitable hand-crafted regularization terms.
With the widespread application of DNNs in medical image processing, numerous DL-based methods have been developed for MAR. [8815915] employed generative adversarial network to correct metal-corrupted sinogram. Since sinogram domain methods are difficult to produce perfect estimation of the metal trace region, secondary artifacts will inevitably be introduced to the reconstructed image. Image domain methods regard the MAR task as image restoration. [8331163] utilized DNNs to generate the prior image that assists in correcting the metal-corrupted regions of the sinogram. [8788607] introduced an unsupervised artifact disentanglement network that produces artifact-reduced images using unpaired training data. The primary limitation of image domain methods is the neglect of the consistency constraint. To address the issue of single domain networks, dual domain methods that exploit dual domain knowledge have been developed these years. The seminal work termed as DuDoNet was proposed by [8953298], and a Radon inversion layer was designed to bridge between the sinogram domain and the image domain network. [9765584] proposed a parallel dual domain network by modifying the fixed priority mode of the previous dual domain network. [10.1007/978-3-030-87231-1_11] proposed a model-driven dual domain network, integrating imaging geometrical model constraints into the mutual learning process of sinogram and image domain information. [WANG2023102729] proposed a deep unfolding dual domain network by unfolding the iterative algorithm for MAR task. Although the dual domain network mentioned above have achieved high-quality CT images, the lack of domain knowledge during training remains a significant limitation in improving the MAR performance. [Shi2024MudNetMD] proposed a method that leverages sinogram domain, image domain, artifact domain, and coding domain knowledge to construct a model-driven network for MAR. [Shi2024CouplingMA] proposed a novel dual domain network that couples model- and data-driven networks for MAR. Although these networks have achieved excellent performance in MAR tasks, they are designed for full-dose CT and do not address the LDMAR issue. Furthermore, the design of these networks has not considered the incorporation of multi-scale and within-scale information. This leaves room for further improvement.
3 The proposed PDuMSRNet for Dual Domain LDMAR Network
[…]
3.1 Model Design and Optimization